Anne and Nancy use a metal alloy that is 15% copper to make jewelry. How many ounces of a 13% alloy must be mixed with a 19% alloy to form 96 ounces of the desired alloy?
Answer by checkley75(3666)

.15*96=.13x+.19(96-x)
14.4=.13x+18.24-.19x
14.4-18.24=.13x-.19x
3.84=.06x
x=3.84/.06
x=64 oz. of 13% alloy is needed.
96-64=32 oz. of 19% alloy is also needed.
proof
.15*96=.13*64+.19*32
14.4=8.32+6.08
14.4=14.4